The resistance of a 2 km length of cable of cross-sectional area 2 mm² and resistivity of

2x10-8 Ωm is


1
Expert's answer
2021-08-23T12:40:27-0400



"R=\rho\cdot\frac{l}{A}=2\cdot10^{-8}\cdot\frac{2000}{2\cdot10^{-6}}=20\ (\Omega)" . Answer